# Deep Unlearn: Benchmarking Machine Unlearning

### v1

@@ 8 IU 0 0 0 4 9 IU 0 0 0 4
Appendix G L2 Distances between model weights
The distance between the Unlearned and Retrained models has also been considered in the literature to evaluate MU. Nevertheless, we observe that models end up at a similar distance to the Retrained model, with significant differences in performance. We further note that one challenging aspect of the L2 distance comparison is the different factors of Weight Decay used by the MU method. The hyper-parameter searches determine these Weight Decay factors, which can significantly vary from one unlearning method to another, making it challenging to compare methods. Furthermore, the best-performing method, MSG, is usually at the same distance as both the Original and Retrained model. For each method, for each initialization seed, we computed the L2 distance between the unlearned model $f_{U}$ and the retrained model $f_{R}$, as well as between the $f_{U}$ and $f_{O}$ (Figure 2).
@@ Appendix G L2 Distances between model weights
Appendix H Requirements

### v2

@@ 8 IU 0 0 0 4 9 IU 0 0 0 4
Appendix C L2 Distances between model weights
The distance between the Unlearned and Retrained models has also been considered in the literature to evaluate MU. Nevertheless, we observe that models end up at a similar distance to the Retrained model, with significant differences in performance. We further note that one challenging aspect of the L2 distance comparison is the different factors of Weight Decay used by the MU method. The hyper-parameter searches determine these Weight Decay factors, which can significantly vary from one unlearning method to another, making it challenging to compare methods. Furthermore, the best-performing method, MSG, is usually at the same distance as the Original and Retained models. For each method, for each initialization seed, we computed the L2 distance between the unlearned model $f_{U}$ and the retrained model $f_{R}$, as well as between the $f_{U}$ and $f_{O}$ (Figure 4).
@@ Appendix C L2 Distances between model weights
Appendix D Data Availability
We will provide the code used to perform unlearning given a trained model, compute the metrics for the hyper-parameter searches, and the evaluation metrics.
Appendix E Requirements
@@
Appendix F Visual Summaries
We provide a visualization of the Performance Retention as they can take the same values.
F.1 Performance Retention
Figure 5: Heatmaps of the performance Retention. (Left) Retain set Retention, (Center) Forget set Retention, (Right) Test Set Retention. (Top) ResNet-18 (Bottom) TinyViT. The optimal value is 1.0, indicating that the model achieves accuracies on par with the reference model.
F.2 Radar Plots
We present Radar Plots illustrating the Retention Deviation using ResNet18 for the methods assessed under U-LIRA.
Figure 6: Radar Plot for Retention Deviation across datasets with the ResNet18 architecture (Fuller is better). Since the Retention Deviation takes values in $[0,+\infty)$, we min-max normalize the values, then subtract them from 1 so that lower values lead to fuller plots.
Figure 7: Radar Plot for the Indiscernibility across datasets with the ResNet18 architecture (Fuller is Better). We first apply a min-max normalization across methods.